OpenStudy (superhelp101):

Could someone please help me with this?:

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

they are typing

OpenStudy (superhelp101):

Ln2-ln(3x+2)=1

OpenStudy (anonymous):

mhmmm

OpenStudy (anonymous):

whats L and N for or are they just normal variables

OpenStudy (superhelp101):

No sorry I meant natural logs

OpenStudy (superhelp101):

I have to get a natural log on both sides of the equation although I forgot how to do that

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

ignore that

OpenStudy (superhelp101):

? do you know how I could get the logs on both sides and btw the answer is 4.125 well at least that is what my teacher got.

OpenStudy (michele_laino):

please note that if I apply the property of the logarithms with same base, I get: \[\ln2-\ln(3x+1)=\ln \left( \frac{ 2 }{ 3x+1 } \right)\] and if I apply the definition of logarithm I get: \[1=\ln e\]

OpenStudy (superhelp101):

I mean 4.214

OpenStudy (superhelp101):

ohhhh I see

OpenStudy (superhelp101):

@Michele_Laino what do I do after applying logs on both sides?

OpenStudy (michele_laino):

you have to equate both numbers, namely, from: \[\ln \left( \frac{ 2 }{ 3x+1 } \right)=\ln e\] we can write: \[\frac{ 2 }{ 3x+1 }=e\] so?

OpenStudy (superhelp101):

would I multiply both of the sides by 3x+1 ?

OpenStudy (superhelp101):

nvm I got how to solve it! thanks for all who came on the question and helped :)

OpenStudy (michele_laino):

Sorry I was at dinner please multiply both sides by 3x+1